Over a period of 180 days, Shawna Johnson earned $450 interest in a retirement account paying interest at a rate of 2%. Find (a) the principal at the beginning of the 180 days and (b) the amount in the account at the end of 180 days. 


(a) The principal was $_______ 
Answer #1

Solution A)

Given: Simple Interest = $450, rate = 2% and time = 180 days = (180/365) years

A)S. I. = ( P * r * t)/100

P* 2 * 3 450= ? 100

P = $45625

B) Amount = Principal + S.I.

Amount = $46075

PS: hit the like button